# Credit Risk Estimation with Non-Financial Features: Evidence from a Synthetic Istanbul Dataset

### v1

@@ Abstract
Financial exclusion constrains entrepreneurship, increases income volatility, and widens wealth gaps. Underbanked consumers in Istanbul often have no bureau file because their earnings and payments flow through informal channels. To study how such borrowers can be evaluated we create a synthetic dataset of one hundred thousand Istanbul residents that reproduces first quarter 2025 TÜİK census marginals and telecom usage patterns. Retrieval augmented generation feeds these public statistics into the OpenAI o3 model, which synthesises realistic yet private records. Each profile contains seven socio demographic variables and nine alternative attributes that describe phone specifications, online shopping rhythm, subscription spend, car ownership, monthly rent, and a credit card flag. To test the impact of the alternative financial data CatBoost, LightGBM, and XGBoost are each trained in two versions. Demo models use only the socio demographic variables; Full models include both socio demographic and alternative attributes. Across five fold stratified validation the alternative block raises area under the curve by about one point three percentage and lifts balanced $F_{1}$ from roughly 0.84 to 0.95, a fourteen percent gain. We contribute an open Istanbul 2025 Q1 synthetic dataset [^1] Code and synthetic dataset: https://github.com/atalaydenknalbant/underbanked_risk_estimation., a fully reproducible modeling pipeline, and empirical evidence that a concise set of behavioural attributes can approach bureau level discrimination power while serving borrowers who lack formal credit records. These findings give lenders and regulators a transparent blueprint for extending fair and safe credit access to the underbanked.
@@ 1 Introduction
Roughly one in three adults worldwide remains “credit invisible,” meaning that mainstream lenders possess too little information to calculate a traditional bureau score (WorldBank2022Findex). This data gap particularly affects workers in the informal economy and newcomers to urban centres who transact largely in cash or through short term instalment plans. Standard scorecards, engineered decades ago for salaried borrowers with multi year repayment histories, reward variables such as revolving balance utilisation, mortgage age, or credit card enquiry rates that underbanked consumers simply do not exhibit. Exclusion from affordable formal credit has measurable macroeconomic consequences; it depresses small business formation, amplifies household income volatility, and entrenches wealth inequality (DemirgucKunt2017InclusionImpact).
In response, both researchers and fintech lenders have turned to alternative data, granular digital traces that originate outside of legacy credit bureaus. Mobile phone metadata record top up regularity, call network diversity, and geo mobility patterns, signals that proxy for income stability and social capital. Utility bill and rent payment histories reveal household budget discipline, while e-commerce receipts show purchasing capacity and behavioral consistency. Smartphone analytics capture device lifecycle length, screen time balance across work and entertainment, and the cadence of ride hailing or food delivery usage; taken together, these metrics form a high frequency portrait of consumption resilience. Short psychometric questionnaires delivered via messaging apps can encode traits such as conscientiousness and risk tolerance that correlate with repayment. When ingested by modern learning pipelines such as gradient boosting ensembles or embedding models, these heterogeneous attributes allow risk scores to be computed for borrowers who lack any conventional repayment history. Early industry deployments report sharper separation of good and bad loans, broader applicant coverage, and lower acquisition costs compared with bureau only approaches (DiMaggio2022InvisiblePrimes)
@@
This paper proposes an open synthetic dataset that mirrors Istanbul 2025 Q1 census and telecom statistics while excluding every form of bureau information. Using that dataset we benchmark off the shelf machine learning models that rely only on alternative signals such as device characteristics, digital commerce activity, and basic socio demographics. Our goal is to quantify the predictive power, robustness, and fairness of bureau free scoring and to provide an extensible test bed for future research on inclusive credit analytics.
@@ 2 Related Work
Research on credit assessment without bureau information has progressed along several complementary lines. One prominent stream leverages mobile phone activity. Analyses of call and text behavior show that network diversity recharge regularity and basic usage statistics can feed logistic and gradient boosting models that separate eventual defaulters from payers even when no financial variables are available (Bjorkegren2020; Oskarsdottir2019BigData). Broader handset analytics that combine call detail records with web browsing categories and social media engagement further raise model stability and invite the use of tree based ensembles and multilayer neural networks (Razavi2025UnlockingCreditAccess). Extension into spatial context demonstrates that adding satellite imagery and public geographic attributes to phone events improves both regression and classification pipelines across logistic forest and boosting families (Simumba2021Spatiotemporal).
@@
Digital interaction footprints gathered at the moment of purchase supply a second evidence track. Data such as device type operating system email domain and transaction timestamp have been placed into logistic regression and random forest frameworks that reach the same predictive quality as commercial bureau scores while reducing reliance on regulated inputs (Berg2020CreditScoringUsingDigitalFootprints). Large consumer platforms report that gradient boosting machines built on in app order cadence spending diversity and session intensity are most informative for young low wealth segments and provide lenders with real time decision capability (Roa2021SuperAppBehavioralPatterns). Complementary experiments using proprietary big data scores generated from millions of behavioral attributes confirm that ensemble learning algorithms based on alternative signals outperform incumbent rule based scorecards designed for bank statements (Jiang2021DecipheringBigData).
@@
Psychometric information offers an orthogonal perspective on borrower reliability. Field pilots in microfinance deploy personality and aptitude questionnaires and compare their output with traditional socio economic scorecards using logistic regression and discriminant analysis. Results show that psychometric variables add measurable lift and that blending them with demographic inputs in hybrid models yields the strongest separation (Sifrain2020PsychometricTesting; Arraiz2015Psychometrics). A similar philosophy underlies studies that draw features from public LinkedIn profiles where stacking ensembles combine demographic career and psycholinguistic indicators to enhance inclusive lending decisions (Alamsyah2025SocialMediaCredit).
@@
Social connectivity also proves valuable. Peer to peer lending datasets augmented with smartphone contact graphs feed random forest and boosting pipelines where sparse networks or strong ties to past defaulters signal elevated risk (Niu2019CreditScoringUsingMachineLearning). Marketplace evidence shows that verified friendships function as an informal screening device lowering interest spreads and observed default within proportional hazards and survival models (Lin2013JudgingBorrowers). Theoretical work on strategic tie formation suggests that when scores incorporate network metrics consumers may consciously restructure links leading to ambiguous changes in overall forecast accuracy (Wei2016CreditScoringWithSocialNetworkData).
@@
Large scale administrative comparisons underscore the practical impact of moving beyond bureau variables. Analyses of consumer lending portfolios reveal that rating grades based on unstructured alternative data preserve their forecasting power over multiple vintages even as their correlation with FICO weakens. Shadow exercises in fintech lending show that logistic models restricted to bureau style inputs would reject substantially more applicants at higher prices whereas gradient boosting systems enriched with behavioural attributes extend affordable credit to many previously excluded borrowers (Jagtiani2019).
@@
Survey and review literature consolidates these findings. A systematic scan of ninety six peer reviewed studies concludes that mobile money digital behaviour and platform based lending constitute the main channels of inclusion and highlights deep learning methods as best suited to heterogeneous inputs (Ha2025). Complementary reviews stress the need for explainable artificial intelligence techniques and propose hybrid supervised unsupervised pipelines that can handle large unstructured alternative datasets while remaining interpretable (Shukla2024; Nwaimo2024).
@@
Although the empirical record is extensive publicly downloadable benchmarks remain scarce. One contribution enriches the Home Credit Default Risk challenge with twenty two engineered behavioural variables and publishes a full evaluation script based on gradient boosting machines yet retains conventional utilisation information (Hlongwane2024). Another releases a cleaned credit score classification corpus together with a CatBoost baseline that relies on occupation age and account derived ratios (Yan2025). All other resources located in the literature either remain proprietary or still contain traditional financial fields. To address this persistent gap the present study generates and publishes a fully synthetic dataset that mirrors Istanbul demographics telecom patterns and digital commerce statistics while excluding every form of bureau information. The corpus supplies only alternative attributes including device characteristics online purchase habits ride hailing frequency and basic socio demographics giving researchers a reproducible foundation for assessing non bureau credit scoring in a major emerging market metropolis.
@@ 3 Methodology
This section explains how the Istanbul 2025 Q1 synthetic dataset was produced, what alternative data features it contains, which learning algorithms were benchmarked, and how discrimination, explainability, fairness, and robustness were evaluated.
@@ 3.1 Synthetic Data Generation
Figure 1 summarises the six stage pipeline that converts public statistics and domain rules into a fully populated record.
Job and income sampling pulls the joint distribution of occupations and pay bands from TÜİK micro tables, then uses weighted random draws to assign a plausible job title and gross salary. Where occupational ambiguity remains, the OpenAI o3 model refines the pick by matching sector keywords that co occur with the chosen education level.
@@
District and rent calculation uses an income ranked map of Istanbul neighbourhoods. Given salary and household size, the algorithm picks an affordable district and computes rent from the district median price per square metre.
Behavioural feature synthesis fills subscriptions, ride hailing intensity, and marketplace activity. Heuristic templates calibrated on industry dashboards provide base rates, and o3 adds small random fluctuations so that behaviour varies naturally across otherwise similar individuals.
@@ 3.2 Feature Catalogue
monthly_subscriptions List List of active digital service subscriptions
@@
Sociodemographics (age, education, home_district) provide essential context for model calibration and fairness checks, ensuring that predictions align with population distributions from TÜİK and enabling the identification of any disparate impacts while also supplying baseline controls that improve overall accuracy by accounting for life stage differences, educational attainment, and regional economic variation.
@@ 3.3 Modeling
Ordered boosting with balanced class weights, Bernoulli row subsampling, and an overfitting detector that stops training after one hundred rounds with no improvement (Dorogush2017CatBoost).
@@
Histogram based tree growth with gradient based one side sampling, balanced class weights, and early stopping after one hundred rounds (Ke2017LightGBM).
@@
Histogram splits, eight tenths row and feature subsampling, balanced loss weighting, and early stopping after one hundred rounds (ChenGuestrin2016XGBoost). Regularisation is set through both $\ell_{1}$ and $\ell_{2}$ penalties.
@@
Elastic net penalty embedded in a preprocessing pipeline that standardises numeric columns and one hot encodes categoricals. The solver runs to full convergence with a tolerance of $10^{-4}$.
@@
Five hundred trees with class balanced bootstrap sampling and a maximum depth of twelve to limit variance.
@@
A single tree with Gini splitting, depth capped at eight, and cost complexity pruning chosen by five fold inner validation.
Hyperparameters for the three boosting libraries are tuned with Bayesian optimisation that uses a tree Parzen estimator for fifty trials (Xia2017BoostedTreeBayes). Random Forest and Decision Tree parameters are selected through grid search, and logistic regression uses cross validated elastic net mixing. All models rely on five fold stratified cross validation so that the prevalence of delinquency_FL is preserved in every fold (Kohavi1995).
Early stopping in the boosters and pruning in the single tree reduce over fitting and keep training times modest. Each model therefore provides a realistic choice for lenders with different hardware budgets and transparency requirements.
@@ 3.4 Alternative Data Impact Study
The incremental value of the behavioural block is assessed by comparing each learner in its Demo variant, which uses only socio demographic inputs,
@@
online_shopping_frequency, social_media_active, monthly_rent}\}.
For each algorithm family the change in area under the receiver operating characteristic curve, precision, recall, and $F_{1}$ is computed. A paired DeLong test evaluates whether the AUC improvement from the Demo to the Full specification is statistically significant in each case (DeLong1988ROC).
Explanatory insight comes from DICE, which produces diverse counter factuals for random validation records (Mothilal2020ExplainingMachine). We tally how often each alternative attribute appears in the minimal edits that flip a prediction. The aggregated counts form an importance profile that pinpoints which phone, commerce, or asset signals contribute most to the uplift observed when moving from Demo to Full models.
@@ 3.5 Evaluation Metrics and Explanation Audits
Discrimination is measured with the area under the curve, precision, recall, and $F_{1}$ are reported at the Youden optimal threshold. Confidence intervals come from bootstrap resamples that draw both rows and feature columns (Efron1979Bootstrap).
@@ 4 Results
Table 2: Five fold out of fold metrics for all models on the synthetic Istanbul dataset. Demo columns use only sociodemographic variables; Full columns add the nine alternative attributes.
@@
Table 2 show results after training all models. The full feature versions of CatBoost, LightGBM, and XGBoost all reach an AUC near 0.965 and an $F_{1}$ close to 0.95, well above the scores of their Demo counterparts. Even the single tree and forest learners gain three to seven AUC points when the alternative block is included, indicating that the uplift stems from the data rather than the modelling technique.

### v3

@@ Abstract
Financial exclusion constrains entrepreneurship, increases income volatility, and widens wealth gaps. Underbanked consumers in Istanbul often have no bureau file because their earnings and payments flow through informal channels. To study how such borrowers can be evaluated we create a synthetic dataset of one hundred thousand Istanbul residents that reproduces first quarter 2025 TÜİK (TURKSTAT) census marginals and telecom usage patterns. Retrieval augmented generation feeds these public statistics into the OpenAI o3 model, which synthesizes realistic yet private records. Each profile contains seven socio demographic variables and ten alternative attributes that describe phone specifications, online shopping rhythm, subscription spend, car ownership, monthly rent, and a credit card flag. To test the impact of the alternative financial data CatBoost, LightGBM, and XGBoost are each trained in two versions. Demo models use only the socio demographic variables; Full models include both socio demographic and alternative attributes. Across five fold stratified validation, the alternative block raises area under the ROC curve by about 1.3 percentage points and lifts balanced $F_{1}$ from roughly 0.84 to 0.95, a 14 percent gain. We contribute an open Istanbul 2025 Q1 synthetic dataset  [^1] Code and synthetic dataset: https://github.com/atalaydenknalbant/underbanked_risk_estimation., a modeling pipeline, and evidence that a concise set of behavioral attributes can approach bureau level discrimination while serving borrowers who lack formal credit records. These findings give lenders and regulators a transparent blueprint for extending safe credit access to the underbanked.
@@ 1 Introduction
Roughly one in three adults worldwide remains “credit invisible” meaning that mainstream lenders possess too little information to calculate a traditional bureau score (Demirgüç-Kunt et al., 2022). This data gap particularly affects workers in the informal economy and newcomers to urban centers who transact largely in cash or through short term installment plans. Standard scorecards, engineered decades ago for salaried borrowers with multi year repayment histories, reward variables such as revolving balance utilization, mortgage age, or credit card inquiry rates that underbanked consumers simply do not exhibit. Exclusion from affordable formal credit has measurable macroeconomic consequences; it depresses small business formation, amplifies household income volatility, and entrenches wealth inequality (Demirgüç-Kunt et al., 2017).
In response, both researchers and fintech lenders have turned to alternative data, granular digital traces that originate outside of legacy credit bureaus. Mobile phone metadata record top up regularity, call network diversity, and geo mobility patterns, signals that proxy for income stability and social capital. Utility bill and rent payment histories reveal household budget discipline, while e-commerce receipts show purchasing capacity and behavioral consistency. Smartphone analytics capture device lifecycle length, screen time balance across work and entertainment, and the cadence of ride hailing or food delivery usage; taken together, these metrics form a high frequency portrait of consumption resilience. Short psychometric questionnaires delivered via messaging apps can encode traits such as conscientiousness and risk tolerance that correlate with repayment. When ingested by modern learning pipelines such as gradient boosting ensembles or embedding models, these heterogeneous attributes allow risk scores to be computed for borrowers who lack any conventional repayment history. Early industry deployments report sharper separation of good and bad loans, broader applicant coverage, and lower acquisition costs compared with bureau only approaches (Maggio et al., 2022)
@@
This paper proposes an open synthetic dataset that mirrors Istanbul 2025 Q1 census and telecom statistics while excluding every form of bureau information. Using that dataset we benchmark off the shelf machine learning models that rely only on alternative signals such as device characteristics, digital commerce activity, and basic socio demographics. Our goal is to quantify the predictive power of bureau free scoring and to provide an extensible test bed for future research on inclusive credit analytics.
@@ 2 Related Work
Research on credit assessment without bureau information has progressed along several complementary lines. One prominent stream leverages mobile phone activity. Analyses of call and text behavior show that network diversity recharge regularity and basic usage statistics can feed logistic and gradient boosting models that separate eventual defaulters from payers even when no financial variables are available (Björkegren and Grissen, 2020; Óskarsdóttir et al., 2019). Broader handset analytics that combine call detail records with web browsing categories and social media engagement further raise model stability and invite the use of tree based ensembles and multilayer neural networks (Razavi and Elbahnasawy, 2025). Extension into spatial context demonstrates that adding satellite imagery and public geographic attributes to phone events improves both regression and classification pipelines across logistic forest and boosting families (Simumba et al., 2021).
@@
Digital interaction footprints gathered at the moment of purchase supply a second evidence track. Data such as device type operating system email domain and transaction timestamp have been placed into logistic regression and random forest frameworks that reach the same predictive quality as commercial bureau scores while reducing reliance on regulated inputs (Berg et al., 2020). Large consumer platforms report that gradient boosting machines built on in app order cadence spending diversity and session intensity are most informative for young low wealth segments and provide lenders with real time decision capability (Roa et al., 2021). Complementary experiments using proprietary big data scores generated from millions of behavioral attributes confirm that ensemble learning algorithms based on alternative signals outperform incumbent rule based scorecards designed for bank statements (Jiang et al., 2021).
@@
Psychometric information offers an orthogonal perspective on borrower reliability. Field pilots in microfinance deploy personality and aptitude questionnaires and compare their output with traditional socio economic scorecards using logistic regression and discriminant analysis. Results show that psychometric variables add measurable lift and that blending them with demographic inputs in hybrid models yields the strongest separation (Sifrain, 2020; Arráiz et al., 2015). A similar philosophy underlies studies that draw features from public LinkedIn profiles where stacking ensembles combine demographic career and psycholinguistic indicators to enhance inclusive lending decisions (Alamsyah et al., 2025).
@@
Social connectivity also proves valuable. Peer to peer lending datasets augmented with smartphone contact graphs feed random forest and boosting pipelines where sparse networks or strong ties to past defaulters signal elevated risk (Niu et al., 2019). Marketplace evidence shows that verified friendships function as an informal screening device lowering interest spreads and observed default within proportional hazards and survival models (Lin et al., 2013). Theoretical work on strategic tie formation suggests that when scores incorporate network metrics consumers may consciously restructure links leading to ambiguous changes in overall forecast accuracy (Wei et al., 2016).
@@
Large scale administrative comparisons underscore the practical impact of moving beyond bureau variables. Analyses of consumer lending portfolios reveal that rating grades based on unstructured alternative data preserve their forecasting power over multiple vintages even as their correlation with FICO weakens. Shadow exercises in fintech lending show that logistic models restricted to bureau style inputs would reject substantially more applicants at higher prices whereas gradient boosting systems enriched with behavioral attributes extend affordable credit to many previously excluded borrowers (Jagtiani and Lemieux, 2019).
@@
Survey and review literature consolidates these findings. A systematic scan of ninety six peer reviewed studies concludes that mobile money digital behavior and platform based lending constitute the main channels of inclusion and highlights deep learning methods as best suited to heterogeneous inputs (Ha et al., 2025). Complementary reviews stress the need for explainable artificial intelligence techniques and propose hybrid supervised unsupervised pipelines that can handle large unstructured alternative datasets while remaining interpretable (Shukla and Gupta, 2024; Nwaimo et al., 2024).
@@
Although the empirical record is extensive publicly downloadable benchmarks remain scarce. One contribution enriches the Home Credit Default Risk challenge with twenty two engineered behavioral variables and publishes a full evaluation script based on gradient boosting machines yet retains conventional utilization information (Hlongwane et al., 2024). Another releases a cleaned credit score classification corpus together with a CatBoost baseline that relies on occupation age and account derived ratios (Yan et al., 2025). All other resources located in the literature either remain proprietary or still contain traditional financial fields. To address this persistent gap the present study generates and publishes a fully synthetic dataset that mirrors Istanbul demographics telecom patterns and digital commerce statistics while excluding every form of bureau information. The corpus supplies only alternative attributes including device characteristics online purchase habits ride hailing frequency and basic socio demographics giving researchers a reproducible foundation for assessing non bureau credit scoring in a major emerging market metropolis.
@@ 3 Methodology
This section explains how the Istanbul 2025 Q1 synthetic dataset was produced, what alternative data features it contains, which learning algorithms were benchmarked, and how discrimination, and explainability were evaluated.
@@ 3.1 Synthetic Data Generation
Figure 1 summarizes the six stage pipeline that converts public statistics and domain rules into a fully populated record.
Job and income sampling pulls the joint distribution of occupations and pay bands from TÜİK (TURKSTAT) micro tables, then uses weighted random draws to assign a plausible job title and gross salary. Where occupational ambiguity remains, the OpenAI o3 model refines the pick by matching sector keywords that co occur with the chosen education level.
@@
District and rent calculation uses an income ranked map of Istanbul neighborhoods. Given salary and household size, the algorithm picks an affordable district and computes rent from the district median price per square meter.
Behavioral feature synthesis fills subscriptions, ride hailing intensity, and marketplace activity. Heuristic templates calibrated on industry dashboards provide base rates, and o3 adds small random fluctuations so that behavior varies naturally across otherwise similar individuals.
@@ 3.2 Feature Catalog
monthly_subscriptions Numeric Total amount paid to active digital service subscriptions
@@
Sociodemographics (age, education, home_district) provide essential context for model calibration checks, ensuring that predictions align with population distributions from TÜİK (TURKSTAT) and enabling the identification of any disparate impacts while also supplying baseline controls that improve overall accuracy by accounting for life stage differences, educational attainment, and regional economic variation.
@@ 3.3 Modeling
Ordered boosting with balanced class weights, Bernoulli row subsampling, and an overfitting detector that stops training after 100 rounds with no improvement (Dorogush et al., 2017).
@@
Histogram based tree growth with gradient based one side sampling, balanced class weights, and early stopping after 100 rounds (Ke et al., 2017).
@@
Histogram splits, eight tenths row and feature subsampling, balanced loss weighting, and early stopping after 100 rounds (Chen and Guestrin, 2016). Regularization is set through both $\ell_{1}$ and $\ell_{2}$ penalties.
@@
Elastic net penalty embedded in a preprocessing pipeline that standardizes numeric columns and one hot encodes categoricals. The solver runs to full convergence with a tolerance of $10^{-4}$.
@@
500 trees with class balanced bootstrap sampling and a maximum depth of 12 to limit variance.
@@
A single tree with Gini splitting, depth capped at 8, and cost complexity pruning chosen by five fold inner validation.
Hyperparameters for the three boosting libraries are tuned with Bayesian optimization that uses a tree Parzen estimator for fifty trials (Xia et al., 2017). Random Forest and Decision Tree parameters are selected through grid search, and logistic regression uses cross validated elastic net mixing. All models rely on five fold stratified cross validation so that the prevalence of delinquency_FL is preserved in every fold (Kohavi, 1995).
Hyperparameter selection is nested within the outer cross validation. For each outer fold, we tune on the training partition only using an inner split and select the configuration that maximizes mean validation AUC. We then refit the model with the selected configuration on the full outer fold training set and evaluate once on the held out fold. This ensures that the held out fold is never used in model selection.
@@ 3.4 Alternative Data Impact Study
The incremental value of the behavioral block is assessed by comparing each learner in its Demo variant, which uses only socio demographic inputs,
@@
For each algorithm family the change in area under the receiver operating characteristic curve, precision, recall, and $F_{1}$ is computed. To test whether adding alternative attributes changes AUC, we generate paired out of fold predictions for the Demo and Full variants under the same five fold split. Each model produces a predicted probability for every record from the fold where that record was held out. We then apply the paired DeLong test to the full set of N paired predictions to assess whether the AUC difference between Demo and Full is statistically significant for each algorithm family (DeLong et al., 1988).
Explanatory insight comes from DICE, which produces diverse counter factuals for random validation records (Mothilal et al., 2020). We tally how often each alternative attribute appears in the minimal edits that flip a prediction. The aggregated counts form an importance profile that pinpoints which phone, commerce, or asset signals contribute most to the uplift observed when moving from Demo to Full models.
@@ 3.5 Evaluation Metrics and Explanation Audits
Discrimination is summarized by the area under the receiver operating characteristic curve. Point metrics, including precision, recall, and $F_{1}$, depend on a decision threshold, so thresholds are selected inside the cross validation loop using the training portion of each fold only. For each fold, we select one operating point on the training fold predictions, the threshold that maximizes $F_{1}$. We then apply the selected threshold to the held out fold and aggregate the resulting out of fold precision, recall, and $F_{1}$ across folds.
@@ 4 Results
Table 2: Five fold out of fold metrics for all models on the synthetic Istanbul dataset. Demo columns use only sociodemographic variables; Full columns add the ten alternative attributes.
@@
Table 2 show the results after training all models. The full feature versions of CatBoost, LightGBM, and XGBoost all reach an AUC near 0.965 and an $F_{1}$ close to 0.95, well above the scores of their Demo counterparts. Even the single tree and forest learners gain three to seven AUC points when the alternative block is included, indicating that the uplift stems from the data rather than the modeling technique.